2012 Nov 06 UT LBC/MODS/LUCI checkout, SXM3 troubleshoot

DThompson (@LBT), MEdwards (@Tucson)

Clear at sunset

M3 troubleshoot

SX M3 was giving problems for LBTI. Checking w/LUCI during twilight. Steve's nightlog will have the details. LUCI images #0001-0014. Switched to MODS to see if it was M2 with problems (seems yes). Restarted OSS server, re-homed M2 hexapod and the source now shows up near the middle of the MODS image (out of focus and with a little coma).

Seems the power supply is dead. We have a spare and James and Dan will install it now.

Unable to get M3 working - no LUCI checkout possible tonight other than internal checks.

LBC

12:45 Found LBC on. Not sure who turned it on…turned it off and back off again

12:50 Typed date with both lbccontrol and obs2. Looks like all is in sync

01:05 Ran 10Bias_Bino - other work was going on so telescope may have been moving slightly, but biases are about 500 counts which is average and the files are writing to /newdata correctly.

01:10 Switching over to LBC

02:24 Slewing to ACT0259

02:34 Running dofpia on ACT0259. Converged. Red is reporting at 0.67"; blue is reporting at 1.3"; DIMM is at 0.8". Looks like blue has a bit of a bad reading, but we'll check again. Blue looks good from the dofpia images

02:46 Running dofpia again; still looks good

02:48 Playing OB for ACT0259 for lbcrangebal check

03:02 lbcrangebal; some small problems with ds9 permissions - solved by closing all ds9 windows and restarting IRAF. Worked fine. IE = -12.39; IA = 4.22

03:08 Slewing to target for evaluation of collimation, check of dithering, We see jumping on the images - both red and blue:

lbcb.20121106.031240 bad; lbcr.20121106.031229 elongated slightly; lbcr.20121106.031622 looks bad; lbcb.20121106.031629 bad....etc.

03:34 Try OB again without guiding. Now images look ok -- Seems to be the same issues as before. John contacts Michele to check for issues

04:00 -- PCS reboot

04:07 Retry with guiding: lbcr.20121106.040254.fits - still bad

04:17 Retry with PA = 90; lbcb.20121106.041715 lbcr.20121106.041709. Still issues. Changing to MODS.

09:18 Switching to LBC to check Michele's patch for guiding

Made TEST2_2012D.ob (in /home/LBTO/dthompson) to point to 7:00 +16:00 (near WT10_269 so we can use that for focus & co-point)

Restarted PCS to update to Michele's patch

DOFPIA at WT10_269

09:16 Running TEST2_2012D.ob ... guiding problem persists. Images around 09:41 actually look OK. On last images of sequence (near 09:45) blue guiding says failed on the UI while red says working. Both images have jumps. Red must be sending offsets that are getting applied to the mount rather than the optics otherwise blue would not have trailed.

09:51 Running TEST2_2012D.ob with guiding off ... images nice and round. Guiding is unchecked on UI but the blue and red channels both say that guiding is working with 8.0s exposures. I don't see a way to take unguided images with one side while the other takes guided images. Unguided red image has 0.56" FWHM (2.5 pixels)!

So...LBC is only usable unguided.

10:13 Switch back to MODS...LBC disconnected from LBT and "other systems" powered down, HK left on.

MODS mods1[r/b].20121106.NNNN

  • lbttcs needed to be restarted
  • modsWake, sieveSnap (#0001s), imSnap (#0002s) look normal
  • data going where it is supposed to
  • 58GB free on /lhome/data

01:45 Preset to check M2 status for JHill and DAshby
r0003-0006 1 sec r_sdss to check M2 status

Starting with AlignCheck scripts, field at high EL
06:25
  • dice5_relrd_02+31.img B=0003 R=0007 to start
  • poor DIQ in the 30s MODS exposures, but FWHM ~0.7", EL~87
  • DIQ noticeably worse on blue images
  • Guide star a bit bright, AGw running with 0.5s exposures

06:39
  • dice5_relxy_02+31.img B=0008 R=0013 to start
  • DIQ better, 0.52" in red, 0.55 in blue.

06:51
  • dice5_xy_02+31.img B=0013 R=0018 EL~83.5@start
  • DIQ still not nice and round over full PSF

07:04
  • dice5_rd_02+31.img B=0018 R=0023 EL~81@start
  • DIQ seems better, still somewhat variable
  • Note: MOVE Type does not appear to update on the GUI (reads Relative when ABS offsets are being done)

Stray Light
07:16
  • SL0305dual_colors.img B=0023 R=0028 EL~63 at start
  • Reflection seen in blue and red, wider (x2 or x3) in blue
  • 2.5 mag star 2 degrees away from this field.

07:34
  • SL0305dual_cardoff.img B=0027 R=0032 EL~63 at start
  • Really really big offsets seem to work fine (>9000 arcsec!). Be interesting to see if we landed where we would expect.

SpecPhot
07:51
  • g191b2b.acq well centered in wide slit after 1 offset.
  • g191b2b.obs will run full script to take all 6 configs.
  • Night seems photometric in at least a qualtitative sense
  • last images B=0042 R=0049, EL~68

10:13 Switching back to MODS.

10:xx Take another set of SpecPhot data on gd71
  • gd71.acq well centered in wide slit after 1 offset.
  • g71.obs will run full script to take all 6 configs.
  • last images B=0054 R=0065, EL~60

12:05 Done with tests I can do without LUCI. Closing early to take calibrations.

Calibrations:
  • 12:17 grpixflats.cal
  • 12:37 grpixflats_r.cal
  • 12:47 grpixflats_b.cal
  • bias8K.cal
  • 13:17 grlamps.cal
  • 13:26 grlamps_r.cal
  • 13:32 grlamps_b.cal

MODS seems to be fine but there is a concern that the drifting that LBTI saw is somehow connected to M2.

LUCI

luci.20121106.NNNN
  • Time checked vs. gps0 (69usec diff)
  • Mask table agrees w/GUI
  • SW NOT restarted
  • No sieve mask images taken
  • Calibration unit and lamps not checked
  • showSupervisorStatus OK
  • ROE/MCE are powered
  • Calib unit is OUT, consistent w/Web-IO
  • /data/luci cleaned up (35GB free)
  • TMGUI updating
  • Dark looked ok, readnoise data taken at end of night
    • MER: 0015 0016 (7.7e-)
    • DCR: 0017 0018 (12.5e-)
  • Data are moving to archive

Unable to check LUCI on sky...unable to control M3 as needed to get things in position.

-- DavidThompson - 06 Nov 2012
Topic revision: r10 - 06 Nov 2012, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback